Write the following cubes in expanded form:
(i) (h • 1, (ii) (2a _ 8bl (iii) [%%+ 1r(iv) [%-!yr.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Ci) (2x + 1>3 • (2x)3 + (1)3 + 3(2xX1)(2x + 1) [Using identity: (x + y)3 = r3 + y3 + 3xy(.r + y)] = 8%3 + 1+ t2x2 + & = &3 + J.2x2 + 6x + 1. (ii) (2.a - 3b>3 = <24>3 -(31>)3 - 3(?,c)(Sb)(?a - Sb) [Using identity: (x - y>3 • %3 - y3 - 3.ry(% -y)] = &3 _ 'l:lb3 _ 36a2b + 54ab2. (w) (ix+ lJ = (i.rJ + (1)3 + 3 x ix x 1 (.r+1) [Using identity: (x + y)3 = .r3 + y8 + 3xy(.r + y )] ;:: 27 .r3 + 1 + 'l:1x2 +-9 .r - 'l:1x3 + 'rt x2 8 4 2 8 4 + 2.r + 1. (iv) (.r-iyJ = (x>3 - (iyJ -3(x>(iy)(x-iy ) [Using identity: (x - y)3 :11 x3 -y3 - 3.ry(x - y)] = .r3 - 7Y3 - 2x2y + i.ry2.